# gamecenter,boxed

### v1

@@ Appendix A Proof of Quantized DP Noise
We argue why quantizing the DP noise is acceptable in Figure 2. In essence, because the sensitive data is already quantized, adding quantized noise to it is equivalent to adding the noise first and then performing the quantization, which aligns with the postprocessing lemma in differential privacy.
@@
We begin by observing that any value, say $v$, can be split into two parts upon quantization: the integral part, $\bar{v}$, and the fractional part, $\hat{v}$, such that $\bar{v}\times\mathbf{2^{\delta}}$ is the (rounded) integral part of $v\times\mathbf{2^{\delta}}$ (i.e. it is $\lceil{v\times\mathbf{2^{\delta}}}\rfloor$), and $\hat{v}$ denotes the fractional part that gets lost due to quantization. This notation can similarly be applied to the noise variable, $\eta$.
@@
$\tilde{v}+\tilde{\eta}=\lceil{v\times\mathbf{2^{\delta}}}\rfloor+\lceil{\eta\times\mathbf{2^{\delta}}}\rfloor=\lceil{(\bar{v}+\eta)\times\mathbf{2^{\delta}}}\rfloor$
The right-hand side represents quantization applied as postprocessing to the sum of the sensitive data and the noise, which adheres to the rules of differential privacy.
@@
$\lceil{(\bar{v}+\eta)\times\mathbf{2^{\delta}}}\rfloor=\lceil{\bar{v}\times\mathbf{2^{\delta}}+\bar{\eta}\times\mathbf{2^{\delta}}+\hat{\eta}\times\mathbf{2^{\delta}}}\rfloor$
We can take out $\bar{v}\times\mathbf{2^{\delta}}$ and $\bar{\eta{}}\times\mathbf{2^{\delta}}$ from the rounding operator because they are exact integers (indeed $\bar{v}\times\mathbf{2^{\delta}}$ is $\lceil{v\times\mathbf{2^{\delta}}}\rfloor$). We then obtain:
$\lceil{(\bar{v}+\eta)\times\mathbf{2^{\delta}}}\rfloor=\lceil{v\times\mathbf{2^{\delta}}}\rfloor+\lceil{\eta{}\times\mathbf{2^{\delta}}}\rfloor+\lceil{\hat{\eta}\times\mathbf{2^{\delta}}}\rfloor$
However, $\hat{\eta}\times\mathbf{2^{\delta}}$ will be lost due to quantization, hence:
$\lceil{(\bar{v}+\eta)\times\mathbf{2^{\delta}}}\rfloor=\lceil{v\times\mathbf{2^{\delta}}}\rfloor+\lceil{\eta{}\times\mathbf{2^{\delta}}}\rfloor=\tilde{v}+\tilde{\eta}$
@@ Appendix B Ablation of Parameters
Figure 9. Ablation over different selections of parameters, $\epsilon=1.0$
Figure 10. Ablation over different selections of $\alpha$, $\epsilon=0.1$
@@
In this section, we empirically justify our choice of hyperparameters: $\eta_{min}$, $\eta_{max}$ and the postprocessing method. We choose these parameters to minimize the Normalized Intra-Cluster Variance (NICV) metric over a set of synthetic datasets. The synthetic datasets were generated using the clusterGeneration package (Qiu and Joe, 2023) varying the number of dimensions ($d$) and clusters ($k$) as $(k,d)\in\{2,4,8,16\}^{2}$, resulting in 16 variations. For each variation, we generate 3 different datasets with the degrees of separation $\{0.25,0.5,0.75\}$. The average cluster size is fixed to $C_{avg}=\frac{2048}{k}$ and the cluster sizes are randomly sampled in the range $[0.70\cdot C_{avg},1.30\cdot C_{avg}]$, leading to a total dataset size in the range $[1433,2662]$.
All the ablation experiments were done on two privacy budget settings: $\epsilon=1.0$ and $\epsilon=0.1$. The results are shown in Figures 9 and 10, respectively. For the min and max cluster experiments, we plot the average of all the NICV values for each $k$ as a separate line and the average of all datasets over all variations as the black dotted line, with the 95% confidence interval shaded around it. We observe that generally, a smaller $\alpha_{min}$ and $\alpha_{max}$ lead to better NICV values, especially for datasets with a high number of clusters. This is expected because a smaller $\alpha_{min}$ gives a tighter bound on the sensitivity and so improves the utility. A smaller $\alpha_{max}$ also imposes a tighter maximum cluster size constraint, which intuitively helps with the convergence of the algorithm. However, the tighter the constraints, the harder it is for them to be satisfied. As we observe in the plots, most datasets with $\alpha_{min}$ or $\alpha_{max}$ set to lower than $1.25$ do not have a value because the min-cost flow solver cannot find a valid assignment. Therefore, we choose $\alpha_{min}=1.25$ and $\alpha_{max}=1.25$ as a balance between utility and feasibility. We also observe that the folding method is better in the low privacy setting ($\epsilon=0.1$), and it does not make a significant difference in the high privacy setting ($\epsilon=1.0$). Therefore, we choose the folding method as the default postprocessing method.

### v2

@@ A.3 Proof of Quantized DP Noise
We argue why quantizing the DP noise is acceptable in Figure 1. In essence, because the sensitive data is already quantized, adding quantized noise to it is equivalent to adding the noise first and then performing the quantization, which aligns with the post-processing lemma in differential privacy.
@@
We begin by observing that any value, say $v$, can be split into two parts upon quantization: the integral part, $\bar{v}$, and the fractional part, $\hat{v}$, such that $\bar{v}\times 2^{q}$ is the (rounded) integral part of $v\times 2^{q}$ (i.e. it is $\lceil{v\times 2^{q}}\rfloor$), and $\hat{v}$ denotes the fractional part that gets lost due to quantization. This notation can similarly be applied to the noise variable, $\eta$.
@@
$\tilde{v}+\tilde{\eta}=\lceil{v\times 2^{q}}\rfloor+\lceil{\eta\times 2^{q}}\rfloor=\lceil{(\bar{v}+\eta)\times 2^{q}}\rfloor$
The right-hand side represents quantization applied as post-processing to the sum of the sensitive data and the noise, which adheres to the rules of differential privacy.
@@
$\lceil{(\bar{v}+\eta)\times 2^{q}}\rfloor=\lceil{\bar{v}\times 2^{q}+\bar{\eta}\times 2^{q}+\hat{\eta}\times 2^{q}}\rfloor$
We can take out $\bar{v}\times 2^{q}$ and $\bar{\eta{}}\times 2^{q}$ from the rounding operator because they are exact integers (indeed $\bar{v}\times 2^{q}$ is $\lceil{v\times 2^{q}}\rfloor$). We then obtain:
$\lceil{(\bar{v}+\eta)\times 2^{q}}\rfloor=\lceil{v\times 2^{q}}\rfloor+\lceil{\eta{}\times 2^{q}}\rfloor+\lceil{\hat{\eta}\times 2^{q}}\rfloor$
However, $\hat{\eta}\times 2^{q}$ will be lost due to quantization, hence:
$\lceil{(\bar{v}+\eta)\times 2^{q}}\rfloor=\lceil{v\times 2^{q}}\rfloor+\lceil{\eta{}\times 2^{q}}\rfloor=\tilde{v}+\tilde{\eta}$
@@ Appendix B Ablation of Parameters
Figure 5: Ablation of $\alpha$ and post-processing strategy on NICV for $\epsilon=0.1$.
@@
In this section, we empirically justify our choice of hyperparameter $\alpha$ and the post-processing strategy. We choose these parameters to minimize the Normalized Intra-Cluster Variance (NICV) metric over a set of synthetic datasets. The synthetic datasets were generated using the clusterGeneration package [58] varying the number of dimensions ($d$) and clusters ($k$) as $(k,d)\in\{2,4,8,16\}^{2}$, resulting in 16 variations. For each variation, we generate three different datasets with the degrees of separation $\{0.25,0.5,0.75\}$. The average cluster size is fixed to $C_{avg}=\frac{2048}{k}$ and the cluster sizes are randomly sampled in the range $[0.70\cdot C_{avg},1.30\cdot C_{avg}]$, leading to a total dataset size in the range $[1433,2662]$.
All the ablation experiments were done on $\epsilon=0.1$. The results are shown in Figure 5, where we plot the average NICV over all datasets for different values of $\alpha$ and post-processing strategies. The average is then divided by the NICV of the naive baseline of no post-processing and no radius constraint. We observe that the “folding” post-processing strategy consistently outperforms the “truncation" and “none” strategies. We also observe that the “Step” strategy outperforms the “Constant” strategy and provides most performance improvement for $\alpha\approx 0.8$. We also note that “radius clipping” provides a consistent improvement. Based on these results, we choose $\alpha=0.8$ and the “Step” constraint strategy with a post-processing strategy of “folding" after “radius clipping” is applied.
B.1 Error Analysis of Simple Gaussian Mechanism
We follow a similar analysis to Section 5.4 to derive the parameters used for GLloyd. Since we do not modify the algorithm (only change the noise distribution), we can use Su et al.’s MSE analysis unchanged:
$\mathsf{MSE}\left(\tilde{\mu}^{(t)}\right)\approx\frac{k^{3}}{N^{2}}\left(\mathsf{Var}\left(\gamma^{S}\right)+4\rho^{2}\mathsf{Var}\left(\gamma^{C}\right)\right)$ (28)
Substituting the variance following our noise mechanism (with the domain-based sensitivity) gives:
$\mathsf{MSE}\left(\tilde{\mu}^{(t)}\right)\approx\frac{k^{3}}{N^{2}}\left(dT(\sigma^{S})^{2}+4\rho^{2}T(\sigma^{C})^{2}\right).$ (29)
Minimizing this using Lagrange multipliers such that:
$\sqrt{\left(\frac{1}{\sigma^{R}}\right)^{2}+\left(\frac{1}{\sigma^{C}}\right)^{2}}=\frac{1}{\sigma}$ (30)
gives the following ratio of noise multipliers:
$\sigma^{C}=\sqrt{\frac{\sqrt{d}}{2\rho}}\sigma^{S}$ (31)
Substituting this back into the MSE equation gives:
$\mathsf{MSE}\left(\tilde{\mu}^{(t)}\right)\approx\frac{k^{3}dT\sigma^{2}(2\rho+\sqrt{d})^{2}}{N^{2}}$ (32)
Setting the per iteration MSE to be less than $0.004$ and rearranging for $T$ gives:
$T\leq\frac{N^{2}(0.004)}{k^{3}dT\sigma^{2}(2\rho+\sqrt{d})^{2}}$ (33)
which we also truncate to be in $[2,7]$.
Appendix C Serverless Extensions
In this section, we discuss alternatives to using the service provider. A trivial alternative is that each client could locally noise their centroids before sending them to another client (or set of clients) to aggregate. The downside to this approach is that the error will compound with the number of clients and decrease utility. An alternative is for all or a subset of the clients to take the role of a computation node in an MPC protocol. In this case, the centroids would be secret shared and aggregated in MPC, then Laplace noise would need to be added similar to FastLloyd. The noise addition can be implemented by each computation node sampling noise locally and adding it to their shares. Depending on how many nodes are assumed, to be honest, this will necessarily add more noise than FastLloyd to ensure privacy in the presence of one or more nodes colluding. Alternatively, the computation nodes can jointly sample the noise using the protocol of Wu et al. [68]. The disadvantage of this approach is that it will add a large computational overhead. However, it would still be significantly faster than related work such as Mohassel et al. [50], which would incur a similar overhead in addition to their already high computation cost.